# One Policy to Run Them All: an End-to-end Learning Approach to Multi-Embodiment Locomotion

### v3

@@ 2 Multi-embodiment Locomotion with a Single Policy
To handle observations of any morphology, URMA first splits the observation vector ${o}$ into robot-specific and general observations ${o_{g}}$, where the former can be of varying size, and the latter has a fixed dimensionality. For locomotion, we subdivide the robot-specific observations into joint and feet-specific observations. This split is not necessary but makes the application to locomotion cleaner. In the following text, we describe everything w.r.t. the joint-specific observations, but the same applies to the feet-specific ones as well. Every joint of a robot is composed of joint-specific observations ${o_{j}}$ and a description vector ${d_{j}}$, which is a fixed-size vector that can uniquely describe the joint by using characteristic properties like the joint’s rotation axis, its relative position in the robot, torque and velocity limits, control range, etc. (see Appendix A for more details). The description vectors and joint-specific observations are encoded separately by the Multilayer Perceptrons (MLPs) ${f_{\bm{\phi}}}$ (with latent dimension $L_{d}$) and ${f_{\bm{\psi}}}$ and are then passed through a simple attention head, with a learnable temperature ${\tau}$ and a minimum temperature ${\epsilon}$, to get a single latent vector
@@
with $f\in\mathcal{F}:\mathcal{X}\rightarrow\mathbb{R}^{J}\times\mathcal{X}$, $h\in\mathcal{H}:\mathbb{R}^{J}\times\mathcal{X}\rightarrow\mathbb{R}^{K}\times\mathcal{X}$ and $w\in\mathcal{W}:\mathbb{R}^{K}\times\mathcal{X}\rightarrow\mathbb{R}$ being the encoder, core and decoder networks, and :×$\ell:\mathbb{R}\times\mathbb{R}\rightarrow[0,1]$ the normalized policy optimization loss function. We quantify the performance of the functions $f,h,w$ with the task-averaged risk
@@ Appendix A Characteristics of the Learning Environment and Robot Embodiments
Every robot is controlled with a PD controller at 50Hz, where the target joint velocity is set to 0, and the target joint position is calculated by: $q^{\text{target}}=q^{\text{nominal}}+\sigma_{a}a$, with $q^{\text{nominal}}$ being a nominal joint position, $a$ the action of the policy and $\sigma_{a}$ a scaling factor. We inherit this action scheme from the implementation details from the DRL locomotion literature [2, 7, 16]. The gains $k_{p}$ and $k_{d}$ and the scaling factor $\sigma_{a}$ vary from robot to robot and are noted in Table 1. The observation space is divided into joint-specific, feet-specific, and general observations. The joint-specific observations ${o_{j}}$ consist of the joint position $q$, joint velocity $\dot{q}$, and previous action $a_{t-1}$ for each joint. The feet observations ${o_{f}}$ include the contact flag $p_{f}$ and the time since each foot’s last contact $p_{f}^{T}$. The general observations ${o_{g}}$ are the linear velocity of the trunk/torso $v$ (only for the critic), angular velocity $\omega$, command x, y and yaw velocities $c$ (re-sampled twice per episode on average), orientation of the gravity vector $g$, current height above ground $h$ (only for the critic), PD controller variables $k_{p}$, $k_{d}$ and $\sigma_{a}$, total mass of the robot $m$ and its dimensions: length $L$, width $W$ and height $H$. All observations are concatenated to $o=[{o_{j}},{o_{f}},{o_{g}}]$ and normalized to $[-1,1]$. For URMA, the joint descriptions ${d_{j}}$ consist of a joint-specific part, which is made of the relative 3D position in the nominal joint configuration, rotation axis, number of direct child joints, joint nominal position, torque limit, velocity limit, damping, rotor inertia, stiffness, friction coefficient and min and max control range, and a general robot part, which is made of $k_{p},k_{d},\sigma_{a},m,L,W,H$. Similarly, the feet descriptions ${d_{f}}$ consist of the relative 3D position of the foot in the nominal joint configuration and of the general robot attributes $k_{p},k_{d},\sigma_{a},m,L,W,H$.
@@ Appendix I Ablation on Humanoids for Zero-Shot Transfer and Fine-Tuning
We start the proof of Theorem 1 by bounding the excess risk, i.e. the difference between the risk $\varepsilon_{\text{avg}}(\hat{f},\hat{h},\hat{w})$ where $\hat{f}$, $\hat{h}$, $\hat{w}$ are the minimizes of the empirical risk, and the minimum task averaged-risk $\varepsilon^{*}_{\text{avg}}$. Notice that excess risk represents how much we pay in terms of performance w.r.t. the optimal policy update since we are optimizing our policy using only a finite amount of samples. We decompose the excess risk, as done in D’Eramo et al. [26], as follows:
@@
We finalize the bound on term $A$ from (11) by substituting (16) in (12):
$\varepsilon_{\text{avg}}(\hat{f},\hat{h},\hat{w})$ $-\frac{1}{nM}\sum_{mi}l(\hat{w}(\hat{h}(\hat{f}(X_{mi}))),Y_{mi})$

### v4

@@ 2 Multi-embodiment Locomotion with a Single Policy
To handle observations of any morphology, URMA first splits the observation vector o{o} into robot-specific and general observations $o_{g}$}, where the former can be of varying size, and the latter has a fixed dimensionality. For locomotion, we subdivide the robot-specific observations into joint and feet-specific observations. This split is not necessary but makes the application to locomotion cleaner. In the following text, we describe everything w.r.t. the joint-specific observations, but the same applies to the feet-specific ones as well. Every joint of a robot is composed of joint-specific observations $o_{j}$} and a description vector $d_{j}$}, which is a fixed-size vector that can uniquely describe the joint by using characteristic properties like the joint’s rotation axis, its relative position in the robot, torque and velocity limits, control range, etc. (see Appendix A for more details). The description vectors and joint-specific observations are encoded separately by the Multilayer Perceptrons (MLPs) $f_{\bm{\phi}}$} (with latent dimension $L_{d}$) and $f_{\bm{\psi}}$} and are then passed through a simple attention head, with a learnable temperature $\tau$} and a minimum temperature $\epsilon$}, to get a single latent vector
@@
with $f\in\mathcal{F}:\mathcal{X}\rightarrow\mathbb{R}^{J}\times\mathcal{X}$, $h\in\mathcal{H}:\mathbb{R}^{J}\times\mathcal{X}\rightarrow\mathbb{R}^{K}\times\mathcal{X}$ and $w\in\mathcal{W}:\mathbb{R}^{K}\times\mathcal{X}\rightarrow\mathbb{R}$ being the encoder, core and decoder networks, and $\ell:\mathbb{R}\times\mathbb{R}\rightarrow[0,1]$ the normalized policy optimization loss function. We quantify the performance of the functions $f,h,w$ with the task-averaged risk
@@
$G(\mathcal{Z}(\bar{\mathbf{X}}))=\mathbb{E}\left[\sup_{z\in\mathcal{Z}}\sum_{mki}\gamma_{mki}z_{k}(X_{mi})|X_{mi}\right],$
@@
[…]
$\,c_{1}\dfrac{L(\ell)L(\mathcal{W})L(\mathcal{H})G(\mathcal{F}(\bar{\mathbf{X}}))}{nM}+c_{2}\dfrac{L(\ell)L(\mathcal{W})\sup_{f}\lVert f(\bar{\mathbf{X}})\rVert O(\mathcal{H})}{nM}$
@@ Appendix A Characteristics of the Learning Environment and Robot Embodiments
Every robot is controlled with a PD controller at 50Hz, where the target joint velocity is set to 0, and the target joint position is calculated by: $q^{\text{target}}=q^{\text{nominal}}+\sigma_{a}a$, with $q^{\text{nominal}}$ being a nominal joint position, $a$ the action of the policy and $\sigma_{a}$ a scaling factor. We inherit this action scheme from the implementation details from the DRL locomotion literature [2, 7, 16]. The gains $k_{p}$ and $k_{d}$ and the scaling factor $\sigma_{a}$ vary from robot to robot and are noted in Table 1. The observation space is divided into joint-specific, feet-specific, and general observations. The joint-specific observations $o_{j}$} consist of the joint position $q$, joint velocity $\dot{q}$, and previous action $a_{t-1}$ for each joint. The feet observations $o_{f}$} include the contact flag $p_{f}$ and the time since each foot’s last contact $p_{f}^{T}$. The general observations $o_{g}$} are the linear velocity of the trunk/torso $v$ (only for the critic), angular velocity $\omega$, command x, y and yaw velocities $c$ (re-sampled twice per episode on average), orientation of the gravity vector $g$, current height above ground $h$ (only for the critic), PD controller variables $k_{p}$, $k_{d}$ and $\sigma_{a}$, total mass of the robot $m$ and its dimensions: length $L$, width $W$ and height $H$. All observations are concatenated to $o=[{o_{j}},{o_{f}},{o_{g}}]$ and normalized to [1,1][-1,1]. For URMA, the joint descriptions $d_{j}$} consist of a joint-specific part, which is made of the relative 3D position in the nominal joint configuration, rotation axis, number of direct child joints, joint nominal position, torque limit, velocity limit, damping, rotor inertia, stiffness, friction coefficient and min and max control range, and a general robot part, which is made of $k_{p},k_{d},\sigma_{a},m,L,W,H$. Similarly, the feet descriptions $d_{f}$} consist of the relative 3D position of the foot in the nominal joint configuration and of the general robot attributes $k_{p},k_{d},\sigma_{a},m,L,W,H$.
@@ Appendix I Ablation on Humanoids for Zero-Shot Transfer and Fine-Tuning
$\mathcal{L}(w(h(f(X_{mi}))),Y_{mi})=-\min\left\{\frac{\pi(a|s,d)}{\pi_{0}(a|s,d)}A^{\pi},\text{clip}(\frac{\pi(a|s,d)}{\pi_{0}(a|s,d)},1-\epsilon,1+\epsilon)A^{\pi}\right\},$
@@
[…]
@@
We finalize the bound on term AA from (11) by substituting (16) in (12):
[…]